A company purchased property for $100,000. The property included a building, a parking lot, and land. The building was appraised at $53,000; the land at $52,200, and the parking lot at $19,800. Land should be recorded in the accounting records with an allocated cost of
Answer #1

Total appraisal cost = 53,000 + 52,200 + 19,800 = 125,000

Amount to be recorded to land

= (52,200/125,000) * 100,000

= 41,760

Comment if you face any issues